OpenStudy (anonymous):

how would I graph -3x - y = -10

OpenStudy (gorv):

first convert it in y=mx+c form

OpenStudy (gorv):

-3x+10=y

OpenStudy (gorv):

y=-3x+10 now for different values of x find y like x=0 than t=10 and plot them

OpenStudy (aum):

To graph a straight line all you need is two points. Put x = 0 and find the y value. That is one point on the y-axis. Put y = 0 and find the x value. That is one point on the x-axis. Using a ruler draw a line passing through those two points.

OpenStudy (mathmale):

Welcome to OpenStudy! Yes, you could first solve the given equation for y, plot the y-intercept (0,10), and then use the slope (m = - 3) to find another point on the line. Or, you could use the concept of vertical and horiz intercepts. Let x=0 and find y. Write the point as (x,y). Next, let y=0 and find x. Write the point as (x,y). Plot the two points and draw a straight line thru them.

OpenStudy (anonymous):

I got (3,-1).. Or would it be (3,1)?

OpenStudy (mathmale):

Where did (3,1) come from? You were asked to graph the line thru the 2 given points, not to solve the equation.

OpenStudy (anonymous):

In my answers on my homework it gives the graphs with the points. two of them didnt make sense and so I went through every step just given to me and I always mix up my negatives and positives... :/

OpenStudy (mathmale):

So, Savvy, what is your question? What kind of help do you need?

OpenStudy (anonymous):

I have to graph them and see where the points intercept to find the solution. I got all mixed up..

OpenStudy (anonymous):

It says "solve the system by graphing" and it gave me: -3x - y = -10 and 4x - 4y = 8 This is my weekest subject :(

OpenStudy (aum):

Plot the first line, -3x - y = -10 using the method in my first reply. Put x = 0: -3(0) - y = -10 0 - y = -10 y = 10 So (0, 10) is one point. Put y = 0: -3x - 0 = -10 x = 10/3. (10/3, 0) is another point Plot these two points and draw a straight line through them. Do a similar procedure with the other line. Find where they intersect and read the x and y coordinates. That will be the answer.

OpenStudy (anonymous):

Thank you so much!! (:

OpenStudy (aum):

You are welcome.